OpenStudy (anonymous):

solving an equation If f(x)=x^3-8x+10, show that there are values c for which f(c) equals; a)pi b)-square root of 3 c)5,000,000

OpenStudy (anonymous):

a. 15.87

OpenStudy (harsimran_hs4):

for each part proceed this way... a) pi let there be a number c such that f(c) = pi i.e x^3 - 8x + 10 = pi x^3 -8x +10-pi = 0 all you need to show is that there is at least one real root of this equation and you are done

OpenStudy (anonymous):

so for b is -9.05

OpenStudy (harsimran_hs4):

what -9.05 ??

OpenStudy (anonymous):

oh I plugged something wrong. Is it, 18.66

OpenStudy (harsimran_hs4):

10 + sqrt(3) = 11.732 x^3 -8x +11.732 = 0 for 2nd part

OpenStudy (klimenkov):

\(f(x)\) is continuous, so it can possess any value between \(f(a)\) and \(f(b)\). If you put \(a=-\infty\) and \(b=\infty\), you will have, that your polynomial can possess any value from \(\mathbb R\). This is the proof.

OpenStudy (anonymous):

Basically, you want to use the intermediate value theorem.